### v2

@@ 4.1 Patient-Event Graph Construction
# EHR_Matrix: Rows represent patients and columns represent medical events
@@
# Calculate posterior probabilities between each pair of events, for the pair of event i and event j:
@@
# Determine edge weights based on the posterior probabilities between each pair of events
edge_weights of (edge between node i and node j) =
posterior probability of (event i and event j)
4.2 Patient-to-Event Graph Learning
In Section 4.1, we mapped each patient sample $\mathbf{X}_{i\cdot}$ to a graph $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, with empirical Bernoulli mean initialization. This transformation converts the original drug recommendation binary classification problem into a graph classification problem. Specifically, we first map the EHR data into a set of vertices and edges $\{0,1\}^{M}\stackrel{{\scriptstyle\text{B.I.}}}{{\rightarrow}}\mathcal{V}_{i},\mathcal{E}_{i}\in[0,1]^{M}$, where B.I. denote Bernoulli Initialization. A graph learning function is then applied to output final drug recommendations $F(\cdot):\mathcal{V}_{i},\mathcal{E}_{i}\rightarrow\{0,1\}^{C}$. We employ EGraphSage [12] as our $F$. We obtain the node and edge embeddings $h_{\nu},h_{e}$ by passing them through a simple MLP. Note that the first layer of the MLP has an interpretation of nonlinearly transforming the Bernoulli statistics residing on the nodes and edges we initialized. The graph learning process aims to discern correlations between events through message aggregation and adjust the event (node) attribute for each patient by updating the node features.
Now for each edge $e_{ij}$ from node $i$ to node $j$, compute the message $m_{j}$ using:
$m_{j}=\text{relu}\left(\mathbf{W}_{m}\times\left(h_{\nu_{j}}\oplus h_{e_{ij}}\right)\right)$
Here, $\oplus$ represents concatenation, $\mathbf{W}_{m}$ is a learnable weight matrix for message passing, and relu is a non-linear activation function.
Message aggregation: After obtaining the message, we aggregate messages $m_{j}$ across edges for each node:
$\text{AGGREGATE}_{i}=\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}m_{j}$
Where $\mathcal{N}(i)$ represents the set of neighbors of the node $i$.
$h_{\nu_{i}}=\text{activation}\left(\mathbf{W}_{u}\times\left(h_{\nu_{i}}\oplus\text{AGGREGATE}_{i}\right)\right),\,\,h_{\nu_{i}}^{\prime}=\frac{h_{\nu_{i}}^{\prime}}{\|h_{\nu_{i}}^{\prime}\|_{2}}$
Where $h_{\nu_{i}}^{\prime}$ is the updated feature of node $i$, and $\mathbf{W}_{u}$ is learnable weight matrix for updating, .$\|.\|_{2}$ denotes the L2 norm.
@@ 4.3 Medication Recommendation
After $K$ iterations, the last layer’s updated node feature $(h^{\prime}_{\nu},\forall{\nu\in\mathcal{V}})$ serves as the output of our graph embedding learning part. After obtaining the node embeddings $h^{\prime}_{\nu}$ for each graph, we concatenate these embeddings to form a single feature vector for each graph. This is represented as: $\mathbf{z}_{\mathcal{G}_{i}}=\oplus_{(\nu\in\mathcal{V})}h^{\prime}_{\nu},$ Where $\mathbf{z}_{\mathcal{G}_{i}}$ is the concatenated feature vector for graph $\mathcal{G}_{i}$ and $\oplus$ denotes the concatenation operation over all nodes (i.e., medical events).
The concatenated feature vectors $\mathbf{z}_{\mathcal{G}_{i}}$ for each graph are then passed through a linear projection for classification. The linear layer has $Y$ independent output dimensions, each corresponding to a binary classification task. Formally, for each output dimension $y$, the operation is:
$\mathbf{o}_{y}=\text{sigmoid}\left(\text{Linear}\left(\mathbf{z}_{\mathcal{G}_{i}}\right)\right)$
$\mathbf{r}_{y}=\begin{cases}1&\text{if }\mathbf{o}_{y}>0.5\\ 0&\text{otherwise}\end{cases}$
Upon processing each graph’s feature vector $\mathbf{z}_{\mathcal{G}_{i}}$, the system yields $Y$ independent outputs. Each of these outputs specifically corresponds to the recommendation decision for a drug. The decision is binary: either the medication is recommended (represented by a value of 1) or not (represented by 0).
@@ 5.1 Databases
We evaluated the performance of BernGraph on two large-scale datasets: the Medical Information Mart for Intensive Care (MIMIC-III) dataset and the Antimicrobial Resistance in Urinary Tract Infections (AMR-UTI). Both databases have been extensively utilized for benchmarking and are available on PhysioNet.
- MIMIC-III represents a vast repository of de-identified health data from intensive care unit (ICU) patients at the Beth Israel Deaconess Medical Center in Boston, Massachusetts [10]. It contains a total of 46520 patients from 2001 to 2012. In this database, the patient data we used was sourced from three sub-datasets, ’ADMISSIONS’, ’DIAGNOSES_ICD’, and ’PROCEDURES_ICD’.
@@
#Medical event 3388 692
@@
- AMR-UTI dataset contains EHR information collected from 51878 patients from 2007 to 2016, with urinary tract infections (UTI) treated at Massachusetts General Hospital and Brigham & Women’s Hospital in Boston, MA, USA [14]. Each patient in this dataset provided a series of binary indicators for whether the patient was undergoing a specific medical event. There are four common antibiotics (medication labels): nitrofurantoin (NIT), TMP-SMX (SXT), ciprofloxacin (CIP) or levofloxacin (LVX).
Data preprocessing. For MIMIC-III, we referred to the data processing released by [36, 33] for fairness. Only the patients with at least 2 visits are incorporated. The medications were selected and retained based on their frequency of occurrence (the top 300). For AMR-URI, we first excluded the basic demographic information including age and ethnicity. The observations that do not have any health event or drug recommendation were removed. For each observation, a feature was constructed from its EHR as a binary indicator for whether the patient is undergoing a particular medical event within a specified time window. After the above preprocessing, we divided both datasets into training, validation and testing by the ratio of 3/5, 1/5, and 1/5. More detailed description of data processing can be found in Appendix 0.B.
5.2 Baselines
We selected three categories of methods as the baseline: statistics-based, RNN-based, and GNN-based models.
- Statistics work. we select the standard Logistic Regression (LR) and Ensemble Classifier Chain (ECC)  [20] as the baseline.
- RNN-based works. RETAIN  [3] incorporates two-level attention with gating mechanisms to detect significant clinical variables. LEAP  [38] is an LSTM-based generation model, leveraging a reinforcement learning module to prevent generating unfavorable drug combinations. DMNC  [11] proposes a new memory-augmented neural network model to improve the patient encoder.
- GNN-based works., GAMENet  [22] integrates the knowledge graph of Drug-Drug Interaction (DDI) by a graph convolutional networks. Further, SafeDrug  [36] combines the drug molecular graph with the DDI graph to predict the safe medication combination. COGNet  [33] retrieves patients’ historical diagnoses and medication recommendations and mines their relationship with current diagnoses via a Conditional Generation Net. In addition to modeling the graph of DDI, MI-GIN [23] directly modeled the EHR data itself in a graph approach. A graph Isomorphism neural network based on multi-task learning is used for medication recommendation.
Moreover, we used the Jaccard Similarity Score (Jaccard), Average F1 (F1), AUROC (Area Under the ROC Curve), Precision-Recall AUC (Area Under the Precision-Recall Curve), and average predicted number of drugs (#Drug) as our evaluation metrics. Each evaluation result was an average of all patients. We further utilized Precision and Recall to evaluate the medication recommendation. The parameter settings can be found in Appendix 0.C.
5.3 Ablation Study
We conducted several analyses to evaluate different modeling methods of node/edge attributes. The main contributions of BernGraph lie in the transformation of binary medical events into continuous values (Bernoulli means). We selected two classic methods for encoding categorical features into continuous values: Log Likelihood Ratio (LLR) and Target Encoding (TE) [21]. The detailed descriptions of these two methods can be found in Appendix 0.D. We further compared against random initialization of nodes and edges for ablation evaluation and assessed BernGraph without node or edge construction. The details of the ablation study are as follows:
[…]
  • BM $w/o$ Post: we maintained the node embedding but replaced the posterior edge embedding with the simple co-occurrence between medical events.
  • LLR $w/$ Post: we replaced the Bernoulli means node embedding with LLR.
  • LLR $w/o$ Post: we used LLR as the simple co-occurrence embedding.
  • TE $w$ Post: we replaced the Bernoulli means node embedding with TE.
  • TE $w/o$ Post: we used TE as the simple co-occurrence embedding.
  • BM $w/$ RE: we replaced the posterior edge embedding with random edge (RE) initialization.
  • RN $w/$ Post: we replaced the Bernoulli means node embedding with random node (RN) initialization.
  • RN $w/$ RE: Both Bernoulli means and posterior embedding were replaced with random initialization.
Moreover, we conducted model ablation to evaluate the event-event graphical correlation modeling compared to the different strategies of simple MLP, LSTM, and Transformer. All of the implementation details of the baseline comparison and ablation study can be found in Appendix 0.E.
Table 2: Performance Comparison on MIMIC-III and AMR-UTI. Best results are highlighted in bold fonts, second-best results are underscored. The results (-) denote models that do not converge, regardless of any model fine-tuning effort.
@@ 5.3 Ablation Study
6 Results
6.1 Comparison Against Baselines
Main Results. Table 2 shows the main results for medication recommendation on two databases. Overall, our proposed BernGraph outperforms all baselines on all metrics (Jaccard, F1, PRAUC, and AUROC). Given that most results were directly collected from the original study [33], our work demonstrates superior performance for the MIMIC-III dataset. It outperforms other approaches, even those that incorporate additional data types, secondary information, and visit modeling for medication recommendations. Some statistical models, such as LR and ECC, outperform early RNN-based baselines LEAP, DMNC, and RETAIN, demonstrating ECC achieves the best results in Jaccard and F1. The difficulties of RNN-based models lie in handling missing historical data, inconsistent data formats in historical records, and the inability to handle situations where no historical data is available. Similar results have been shown in AMR-UTI; given that six visit datasets could serve as the RNN’s timesteps, it often vanishes gradients within various model fine-tunings. Graph-based methods tend to converge well and result in better performances compared to RNN baselines. This shows the effectiveness of modeling event-event correlations and structure information. While BernGraph is solely based on binary medical events, further integrating this binary data modeling with external EHR information, like DDI, could enhance recommendation performance.
Visit-Included Results. The results presented in Table 2 exclude information derived from patient visits. To provide a more comprehensive assessment incorporating visit data, we present an average recommendation performance across all patient visits in Table 3. Specifically, some of the baseline methods rely on sequential learning. In these approaches, during each visit, every medication is allocated a recommendation probability upon its generation.
Model Jaccard F1 PRAUC AUROC #Drug
LR 0.4844 0.6483 0.7492 0.9029 16.1221
ECC 0.5003 0.6590 0.6802 0.9214 18.4127
RETAIN 0.4876 0.6484 0.7594 0.9225 19.8297
LEAP 0.4487 0.6119 0.6480 0.8927 19.1095
DMNC 0.4834 0.6492 0.7569 0.9237 20.1087
GAMENet 0.5084 0.6648 0.7670 0.9235 26.1983
SafeDrug 0.5276 0.6692 0.7649 0.9218 19.8847
COGNet 0.5127 0.6850 0.7739 0.9266 27.3277
MT-GIN 0.5273 0.7527 0.7499 0.9188 15.1233
Proposed 0.5902 0.8465 0.8448 0.9635 15.6944
We first average their performance over multiple visits for each patient and then further average these results across all patients. In this context, BernGraph shows an effective and robust performance and outperforms all baselines.
@@ 6.1 Comparison Against Baselines
Medication Recommendation Analysis. Fig. 3 provides an extensive comparison of Precision and Recall metrics for medication recommendation across all baselines and BernGraph. Hence, we employ several evaluation metrics to ensure a comprehensive comparison of our approach against existing methods. One prevalent issue is the False Positive (FP) problem, where the recommendation system suggests an unsuitable or unnecessary drug to a patient. This results in a relatively high recall but a low precision due to incorrect recommendations. On the other hand, the False Negative (FN) issue happens when the recommendation system fails to suggest a drug that would have been beneficial for the patient. This results in low recall because it fails to capture a relevant recommendation. However, precision can still be high for recommendations. In Fig. 3, the ground truth of the average number of recommendations is 19.2268 in the MIMIC-III dataset BernGraph has shown a similar performance with 15.7129 average recommendations. This indicates our method tends to recommend precise (high Jaccard and Precision scores) but fewer drugs out of 131 while still achieving competitive recall rates.
Table 4: Ablation study on node and edge embeddings. We compare various embedding methods for node/edge attributes to validate the contributions of the proposed embedding approach. Bold: best; underscored: second-best.
@@ 6.2 Ablation Study
For ablations on node and edge embeddings, Table 4 shows that among various embedding strategies on both datasets, combining our Bernoulli mean node embedding with posterior probability edge embedding achieves the best results. If we compare the contributions of node (Bernoulli mean) and edge (posterior probability) embeddings to the graph-based drug recommendation results, we find that node embedding has a relatively more substantial impact. The random node destroys the recommendation performance even when considering posterior probabilities, resulting in a significantly low Jaccard score of 0.29. However, using the Bernoulli mean for edge embedding initialization yields a much better performance, with a Jaccard score of 0.5692. This could be because node embedding directly represents the original EHR data, while edges are inferred based on the original data to model the relationship between different events. On the other hand, the purpose of using GNN is to capture event-event correlations. Therefore, edge embeddings merely aid in GNN computations. The results show that even without initializing edge embeddings with posterior probabilities, GNN can still capture event-event correlations, though with a performance decrease. For model ablation, GNN outperforms all baselines (i.e., MLP, LSTM, and Transformer) on two datasets.
@@
Figure 4: Visualization of features learned by BernGraph. Each node denotes an event, and edges represent event-event correlations. A thicker edge shows a stronger event-event association, and nodes with higher values are highlighted.
@@ 6.3 Visualization of Recommendations
Fig. 4 shows representations learned by GNN for two different patients on a subset of EHR medical events. GNN produces distinct feature values for different nodes, for visualization and better understanding, we processed the node feature values using the TOP-K approach, highlighting the top K nodes with the highest feature values. It can be observed that GNN has learned completely different representations for different patients and has emphasized different features on various nodes. For instance, the "Hyperlipidemia" node, which is highlighted in Patient B’s data, receives a lower weight in Patient A’s data. This indicates that this event is more critical in Patient B’s data compared to Patient A. Moreover, GNN has learned certain connections between important features for each patient. This result highlights how the same event can exhibit entirely different levels of importance across different patients’ data. For example, in Patient B’s data, three identical series (MDII) of medical events are simultaneously highlighted. Through this simple analysis, we confirmed that GNN learned different representations for different patients, providing valuable information for downstream tasks. More results can be found in Appendix 0.F.

### v3

@@ 4.2 Patient-to-Event Graph Learning
# EHR_Matrix: Rows represent patients, and columns represent medical events
@@
# Calculate posterior probabilities between each pair of events for the pair of event i and event j:
@@
# Determine edge weights by the posterior probabilities between each pair of events
edge_weights of edge from node i to node j =
posterior probability of (i, j)
@@ 4.3 Medication Recommendation
After $K$ iterations, the last layer updated node feature $(h^{\prime}_{\nu},\forall{\nu\in\mathcal{V}})$ serves as the output of our graph embedding learning part. After obtaining the node embeddings $h^{\prime}_{\nu}$ for each graph, we concatenate these embeddings to form a single feature vector for each graph. This is represented as:
$\mathbf{z}_{\mathcal{G}_{i}}=\oplus_{(\nu\in\mathcal{V})}h^{\prime}_{\nu},$ (6)
where $\mathbf{z}_{\mathcal{G}_{i}}$ is the concatenated feature vector for graph $\mathcal{G}_{i}$ and $\oplus$ denotes the concatenation operation over all nodes (i.e., medical events). The concatenated feature vectors $\mathbf{z}_{\mathcal{G}_{i}}$ for each graph are then passed through a linear projection for classification. The linear layer has $Y$ independent output dimensions, each corresponding to a binary classification task. Formally, for each output dimension $y$:
$\mathbf{o}_{y}=\text{sigmoid}\left(\text{Linear}\left(\mathbf{z}_{\mathcal{G}_{i}}\right)\right)$ (7)
$\mathbf{r}_{y}=\begin{cases}1&\text{if }\mathbf{o}_{y}>0.5\\ 0&\text{otherwise}\end{cases}$ (8)
Upon processing graph feature vectors $\mathbf{z}_{\mathcal{G}_{i}}$, the system yields $Y$ independent outputs. Each of these outputs corresponds explicitly to the decision to recommend drugs. The decision is binary: either the medication is recommended (represented by a value of 1) or not (denoted by 0).
@@ 5.1 Databases
We evaluated BernGraph on two large-scale datasets: the Medical Information Mart for Intensive Care (MIMIC-III) dataset and the Antimicrobial Resistance in Urinary Tract Infections (AMR-UTI). Both databases have been extensively utilized for benchmarking and are available on PhysioNet.
- MIMIC-III represents a vast repository of de-identified health data from intensive care unit (ICU) patients at the Beth Israel Deaconess Medical Center in Boston, Massachusetts (Johnson et al., 2016). It contains a total of 46520 patients from 2001 to 2012.
- AMR-UTI dataset contains EHR information collected from 51878 patients from 2007 to 2016, with urinary tract infections (UTI) treated at Massachusetts General Hospital and Brigham & Women’s Hospital in Boston, MA, USA (Michael et al., 2020). This dataset provided a series of binary indicators for whether a patient was undergoing a specific medical event. There are four common antibiotics (medication labels): nitrofurantoin (NIT), TMP-SMX (SXT), ciprofloxacin (CIP), or levofloxacin (LVX).
@@
#Patients 46250 51878
Avg. event sequence length 3388 692
@@
Data preprocessing. For MIMIC-III, we referred to the data processing released by (Yang et al., 2021b; Wu et al., 2022) for fairness. Only the patients with at least 2 visits are incorporated. The medications were selected and retained based on their frequency of occurrence (the top 300). For AMR-URI, we first excluded the basic demographic information including age and ethnicity. The observations that do not have any health event or drug recommendation were removed. For each observation, a feature was constructed from its EHR as a binary indicator for whether the patient is undergoing a particular medical event within a specified time window. After the above preprocessing, we divided both datasets into training, validation and testing by the ratio of 3/5, 1/5, and 1/5. More detailed description of data processing can be found in Appendix A.
Table 2: Performance Comparison on MIMIC-III and AMR-UTI. Best results are highlighted in bold fonts, second-best results are underscored. The results (-) denote models that do not converge.
@@
5.2 Baselines
- Statistics-based methods: We use standard Logistic Regression (LR) and Ensemble Classifier Chain (ECC) (Read et al., 2009) as baselines. - RNN-based methods: RETAIN (Choi et al., 2016a) applies two-level attention with gating for key clinical variables. LEAP (Zhang et al., 2017) uses an LSTM-based model with reinforcement learning to avoid adverse drug combinations. DMNC (Le et al., 2018) introduces a memory-augmented network to enhance patient encoding. - GNN-based methods: GAMENet (Shang et al., 2019) incorporates Drug-Drug Interaction (DDI) knowledge using graph convolutional networks. SafeDrug (Yang et al., 2021b) combines the drug molecular graph with DDI to predict safe medication combinations. Using a Conditional Generation Net, COGNet (Wu et al., 2022) mines relationships between historical and current diagnoses. MI-GIN (Shu et al., 2023) directly models EHR data using a graph isomorphism neural network for medication recommendation. - Evaluation: We evaluated models using the Jaccard Similarity Score, Average F1, AUROC, Precision-Recall AUC, and average predicted number of drugs (#Drug), averaging results overall for patients. Precision and Recall were also used to assess recommendation performance. Parameter settings are detailed in Appendix B.
@@ 5.2 Baselines
Table 3: An ablation study on embeddings is conducted to validate the contributions of the proposed embedding approach. The best-performings are highlighted in bold, while the second-best results are underscored.
@@ 5.3 Ablation Study
We evaluated different node/edge modeling methods, with our key contribution being the transformation of binary medical events into continuous Bernoulli means. We compared BernGraph with two classic encoding methods, Log Likelihood Ratio (LLR) and Target Encoding (TE) (Rodríguez et al., 2018) (details in Appendix C). Additionally, we performed ablation studies using random node/edge initialization and assessed BernGraph without node or edge construction. We evaluated several variations of our method. BM $w/$ Post is our proposed approach. BM $w/o$ Post uses simple co-occurrence for edge embeddings. LLR $w/$ Post replaces Bernoulli means with LLR for node embeddings, while LLR $w/o$ Post uses LLR for both. TE $w/$ Post and TE $w/o$ Post follow a similar pattern with Target Encoding. BM $w/$ RE uses random edge initialization, and RN $w/$ Post and RN $w/$ RE replace node or both node/edge embeddings with random initialization. We also conducted ablation studies comparing GNN event-event modeling with MLP, LSTM, and Transformer. Details are in Appendix D.
6 Results
6.1 Comparison Against Baselines
Main Results. Table 2 presents the main results for medication recommendation on two datasets. Our proposed BernGraph outperforms all baselines across all metrics (Jaccard, F1, PRAUC, AUROC), particularly excelling on the MIMIC-III dataset, even against models using additional data and visit modeling. Statistical models like LR and ECC surpass early RNN-based methods (LEAP, DMNC, RETAIN), highlighting the strong performance of ECC. RNN models struggle with missing or inconsistent historical data, while graph-based methods perform better by modeling event-event correlations, showing the effectiveness of modeling event-event correlations and structure information. BernGraph, based solely on binary events, could improve by incorporating external EHR data like DDI.
Visit-Included Results. The results presented in Table 2 exclude information derived from patient visits. To provide a more comprehensive assessment incorporating visit data, we present an average recommendation performance across all patient visits in Table 4. Specifically, some of the baseline methods rely on sequential learning. In these approaches, every medication is allocated a recommendation probability upon generation in each visit. We first average their performance over visits for each patient and then further average these results across all patients. In this context, BernGraph shows an effective and robust performance and outperforms all baselines.
Medication Recommendation Analysis. Fig. 3 extensively compares Precision and Recall metrics for medication recommendation. We employ several evaluation metrics to compare BernGraph against existing methods. The False Positive (FP) issue arises when the system suggests unnecessary drugs, leading to high recall but low precision. Conversely, False Negatives (FN) occur when beneficial drugs are missed, resulting in low recall but potentially high precision. As shown in Fig. 3, we provide 15.7129 recommendations on average, and the ground truth for MIMIC-III is 19.2268. This reflects that our method can recommend fewer but more precise drugs, maintaining competitive recall.
Model Jaccard F1 PRAUC AUROC #Drug
LR 0.4844 0.6483 0.7492 0.9029 16.1221
ECC 0.5003 0.6590 0.6802 0.9214 18.4127
RETAIN 0.4876 0.6484 0.7594 0.9225 19.8297
LEAP 0.4487 0.6119 0.6480 0.8927 19.1095
DMNC 0.4834 0.6492 0.7569 0.9237 20.1087
GAMENet 0.5084 0.6648 0.7670 0.9235 26.1983
SafeDrug 0.5276 0.6692 0.7649 0.9218 19.8847
COGNet 0.5127 0.6850 0.7739 0.9266 27.3277
MT-GIN 0.5273 0.7527 0.7499 0.9188 15.1233
Proposed 0.5902 0.8465 0.8448 0.9635 15.6944
6.2 Ablation Study
@@
Table 3 shows that combining Bernoulli mean node embedding with posterior probability edge embedding yields the best results. If we compare the contributions of node (Bernoulli mean) and edge (posterior probability) embeddings to the graph-based drug recommendation results, we find that node embedding has a relatively more substantial impact. The random node destroys the recommendation performance even when considering posterior probabilities, resulting in a low Jaccard score of 0.29. However, using the Bernoulli mean for edge initialization yields a much higher Jaccard score of 0.5692. This could be because node embedding directly represents the original EHR data, while edges are inferred based on the original data to model the relationship between different events. Therefore, edge embeddings merely aid in GNN computations. The results show that even without initializing edge embeddings with posterior probabilities, GNN can still capture event-event correlations, though with a performance decrease. For model ablation, GNN outperforms all baselines.
@@
Figure 4: Feature visualization: Each node denotes an event, and edges represent event-event correlations. A thicker edge shows a stronger event-event association, highlighting higher-value nodes.
@@ 6.3 Visualization of Recommendations
Fig. 4 shows GNN-learned representations for two patients on a subset of EHR events. We highlight the top nodes with the highest feature values by TOP-K, revealing that GNN captures distinct representations for each patient. For example, the "Hyperlipidemia" node is more prominent in Patient B than in Patient A, indicating greater relevance. GNN also learns key feature connections, with certain event series highlighted in Patient B. This demonstrates the ability of GNN to differentiate between patients and provide valuable insights for downstream tasks.